OpenStudy (anonymous):

Suppose \(\lim_{x\rightarrow a}f(x) = c\). How can I show that \(\lim_{x\rightarrow a} \sqrt[n]{f(x)} \) exists for all n in \(\mathbb{N}\)?

OpenStudy (perl):

looks like an inductive proof

OpenStudy (perl):

that doesn't mean its bad :)

OpenStudy (perl):

with n=1 , thats true by limit rules

OpenStudy (freckles):

I guess we are suppose to just assume a and c>0?

OpenStudy (perl):

oh, when n=1 , its not square rooting at all !

OpenStudy (anonymous):

\(a \in \mathbb{R}\) and c>0

OpenStudy (perl):

when n=1, you have lim (f(x)) ^(1/1) = lim f(x)

OpenStudy (anonymous):

A side question- is the limit c^(1/n)?

OpenStudy (perl):

yes

OpenStudy (perl):

because

OpenStudy (perl):

draw button died :/

OpenStudy (perl):

lim (sqrt(f(x)) = sqrt( lim(f(x))

OpenStudy (anonymous):

lim (sqrt(f(x)) = sqrt( lim(f(x)) <- prove it!

OpenStudy (perl):

actually that is the case n=2

OpenStudy (perl):

the statement is more general

OpenStudy (perl):

lim [ f(x)^(1/n)] = ( lim(f(x)) ^(1/n)

OpenStudy (anonymous):

I'll try induction. Have to go for a while, will get back to this soon.

OpenStudy (perl):

lim [ f(x) ^(1n)] = [ lim(f(x) ] ^(1/n)

OpenStudy (perl):

freckles wanted to say something :)

OpenStudy (freckles):

say we have \[f(x)=\frac{c}{a} x \\ \lim_{x \rightarrow a} f(x)=c \\ \lim_{x \rightarrow a}\sqrt{f(x)}=\lim_{x \rightarrow a} \sqrt{ \frac{c x}{a}}= \sqrt{c}\] But I think a can be a real number not equal to 0 for this function since we cannot divide by 0 - i'm just worried about the restrictions is all

OpenStudy (perl):

this question has multiple levels of generality ( i dont know a better word for it)

OpenStudy (perl):

you *found* a special case of n = 2 :)

OpenStudy (perl):

it really stinks you can't edit your comments

OpenStudy (freckles):

i know what you meant though

OpenStudy (freckles):

And I was just picking a function that satisfied c>0 and a is a real number but I think there needs to be more restrictions on a or maybe on f

OpenStudy (freckles):

Maybe I'm being difficult

OpenStudy (perl):

yeah i see what you mean

OpenStudy (perl):

so you would just have to say, a =/= 0

OpenStudy (anonymous):

We shall prove it in a more general form: \(\lim_{x\rightarrow a} \sqrt[n]{f(x)^m}\). First, we'll show that if \(\lim_{x\rightarrow a}f(x) = c\), then \(\lim_{x\rightarrow a}f(x)^m \) exists and \(\lim_{x\rightarrow a}f(x)^m =(\lim_{x\rightarrow a}f(x))^m = c^m\). This shall be proved by induction. When m=2, \(\lim_{x\rightarrow a}f(x)^2 = \lim_{x\rightarrow a}f(x) \cdot \lim_{x\rightarrow a}f(x) =(\lim_{x\rightarrow a}f(x))^2= c^2\) Suppose \(\lim_{x\rightarrow a}f(x)^k = (\lim_{x\rightarrow a}f(x))^k=c^k\) for some positive integers k. Consider m=k+1, \(\lim_{x\rightarrow a}f(x)^{k+1} \) \(= \lim_{x\rightarrow a}f(x)^k \cdot \lim_{x\rightarrow a}f(x) \) \(= (\lim_{x\rightarrow a}f(x))^k \cdot\lim_{x\rightarrow a}f(x) \) \(= c^k\cdot c = c^{k+1}\) Hence, by principle of mathematical induction, if \(\lim_{x\rightarrow a}f(x) = c\), \(\lim_{x\rightarrow a}f(x)^m = (\lim_{x\rightarrow a}f(x))^m =c^m\) for all positive integers m. Now, let \(g(x) = \sqrt[n]{f(x)^m}\), then we have \(g(x)^n = f(x)^m\), and \(\lim_{x\rightarrow a}g(x)^n = \lim_{x\rightarrow a}f(x)^m\) From the above, we've shown that \(\lim_{x\rightarrow a}f(x)^m = c^m\). Hence, we have \((\lim_{x\rightarrow a}g(x))^n = \lim_{x\rightarrow a}g(x)^n = \lim_{x\rightarrow a}f(x)^m = c^m\) That is, \((\lim_{x\rightarrow a}g(x))^n= c^m\)Taking the nth root on both sides, we have \(\lim_{x\rightarrow a}g(x)= c^\frac{m}{n}\) Hence, we've proved that if \(\lim_{x\rightarrow a}f(x) = c\), \(\lim_{x\rightarrow a} \sqrt[n]{f(x)^m}\) exists, and \(\lim_{x\rightarrow a} \sqrt[n]{f(x)^m}= c^\frac{m}{n}\) for all \(m,n\in\mathbb{N}\) Let m=1, we have if \(\lim_{x\rightarrow a}f(x) = c\), then \(\lim_{x\rightarrow a} \sqrt[n]{f(x)}\) exsits, and \(\lim_{x\rightarrow a} \sqrt[n]{f(x)}= c^\frac{1}{n}\) for all \(n\in\mathbb{N}\)